a block of wood is being pushed by a boy .the mass of the block is 50kg if the coefficient of static friction is 0.5 then what should be the minimum force appliedby the boy to push the block of wood?

Grade:10

1 Answers

Sayak Biswas
11 years ago

Ok watch

 

mass of block=50kg

coeff.of static friction=0.5

Tking vertical axis,

N-mg=0(since no motion in vertical direction)

or n=mg

therefore,maximum static friction=Coeff.of friction*Normal reaction 

                                   =0.5*mg

                                    =0.5*50*10
                                     =250 Newtons

 If the boy applies force equal to 250N,the block would tend to move